# New Integrable Chiral Cosmological Models with Two Scalar Fields

### v1

@@ IV.1 The FLRW metric with conformal time
V Potentials and solutions of the field equations
V.1 The way to get the general cosmological solution of the CCM
In the FLRW metric with the conformal time, Eqs. (22)–(24) have the following form:
@@ V.1 The way to get the general cosmological solution of the CCM
Our goal is to get the general solutions of these equations for the potentials given by the formula (21). If we know the general solutions for the corresponding model in the Jordan frame, or, in other words, if the functions $a(\tau)$, $\sigma^{1}(\tau)$, and $\sigma^{2}(\tau)$ are known, then the scale function $a_{\mathrm{E}}$ is given by Eq. (52) and the Hubble parameter
@@
Here we use the fact that conformal times for both frames coincide (up to an additive constant, which can be set to be zero without loss of generality), since $dt/a=dt_{\mathrm{E}}/a_{\mathrm{E}}$. Having these expressions for $a_{\mathrm{E}}$, $H_{\mathrm{E}}$,
@@
It is not always possible to integrate Eq. (37). To integrate field equations in the Jordan frame we must firstly choose a particular form of the potential $V$. Below we present the general solutions for three types of the potential $V$. The obtained solutions allow us to get the general solutions for the corresponding CCM.
V.2 Potential $V=\frac{1}{2}{R_{0}U_{0}}$
@@
$V_{\mathrm{E}}=\frac{M^{4}_{\mathrm{Pl}}R_{0}}{8U_{0}}\cosh^{2}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\cosh^{2}\left(\frac{\chi^{2}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\,.$ (60)
@@
$y^{A}=\left\{\begin{aligned} B_{1}^{A}\sin(\sqrt{K}\tau)+B_{0}^{A}\cos(\sqrt{K}\tau),\quad&\quad K>0;\\ B_{1}^{A}\tau+B_{0}^{A},\qquad\qquad\qquad\quad\quad\quad\,\,\,&\quad K=0;\\ B_{1}^{A}\sinh(\sqrt{K}\tau)+B^{A}_{0}\cosh(\sqrt{K}\tau),&\quad K<0;\\ \end{aligned}\right.$ (61)
@@
V.3 Potential $V=\frac{1}{2}{R_{0}U_{0}}+c_{1}\left(\sigma^{1}\right)^{4}+c_{2}\left(\sigma^{2}\right)^{4}$
@@
For this potential $V$, system (37) is completely separable, and we have 2 independent integrals of motion:
@@
The form of solutions of (63) depends on the signs of $c_{A}$ and $E_{A}$. For $K=0$, we obtain:
@@
    $y^{A}(\tau)=\left(\frac{E_{A}}{c_{A}}\right)^{\frac{1}{4}}\mathrm{cn}\left.\left(\pm 2\left(E_{A}c_{A}\right)^{\frac{1}{4}}(\tau-\tau_{i})+\mathrm{cn}^{-1}\left(u_{n}\left|\frac{1}{2}\right.\right)\right|\frac{1}{2}\right);$ (64)
@@
Here $u_{A}=|c_{A}/E_{A}|^{1/4}y^{A}(\tau_{i})$, and the choice of sign in “$\pm$” is determined by initial conditions. Obviously, there are no solutions when $E_{A}<0$ and $c_{A}>0$.
@@
If we put $\chi^{2}=0$, then $V_{\mathrm{E}}$ coincides with the potential from single-field model proposed by Bars and Chen in Ref. [5]:
@@
So, the proposed integrable model with potential (67) is two-field generalization of the known single-field integrable model [5].
V.4 Potential $V=\frac{1}{2}{R_{0}U_{0}}+c\left(\left(\sigma^{1}\right)^{2}+\left(\sigma^{2}\right)^{2}\right)^{2}$
@@
$V_{2}=\frac{RU_{0}}{2}+c\left(\left(\sigma^{1}\right)^{2}+\left(\sigma^{2}\right)^{2}\right)^{2},$ (69)
where $c$ is a nonzero constant, general solutions can also be explicitly written in terms of the Jacobi elliptic functions. In this case, Eq. (38) with $K=0$ takes the form
$\frac{1}{2}\left(\frac{dy^{1}}{d\tau}\right)^{2}+\frac{1}{2}\left(\frac{dy^{2}}{d\tau}\right)^{2}+c\left(\left(y^{1}\right)^{2}+\left(y^{2}\right)^{2}\right)^{2}=E.$ (70)
By introducing “polar coordinates”,
@@
$\frac{1}{2}\left(\frac{d\rho}{d\tau}\right)^{2}+\frac{1}{2}\rho^{2}\left(\frac{d\theta}{d\tau}\right)^{2}+c\rho^{4}=E.$ (72)
In analogy with classical mechanics, one may observe that $\theta$ is a cyclic coordinate, which means that
@@
$\frac{1}{2}\left(\frac{d\rho}{d\tau}\right)^{2}+\frac{L^{2}}{2\rho^{2}}+c\rho^{4}=E.$ (74)
In what follows, we restrict ourselves to the case of $c>0$ only. In such a case, the solution for $\rho$ reads
@@
$\rho^{2}_{0,1}=2\sqrt{\frac{E}{3c}}v_{0,1},\quad v_{k}=\cos\left(\frac{1}{3}\arccos\left(-\left(\frac{E_{min}}{E}\right)^{3/2}\right)-\frac{2\pi k}{3}\right),\quad k=0,1,2,$
$\beta=\frac{1}{4}\left(\frac{3}{cE}\right)^{1/4},\quad E_{min}=\frac{3}{2}\left(\frac{cL^{4}}{2}\right)^{1/3}.$
Note that since $E\geqslant E_{min}$, we always have $v_{0}\geqslant v_{1}>0>v_{2}$. The equality $v_{0}=v_{1}$ holds when $E=E_{min}$. The solution for $\theta$ is
$\theta(\tau)=\frac{\mathrm{sgn}(L)}{2v_{0}\sqrt{v_{0}-v_{2}}}\left(\frac{E_{min}}{E}\right)^{3/4}\Pi\left(\left.1-\frac{v_{1}}{v_{0}};\frac{\sqrt{v_{0}-v_{2}}}{2\beta}\tau+C_{\rho}\right|\frac{v_{0}-v_{1}}{v_{0}-v_{2}}\right)+C_{\theta}.$ (76)
@@
If $E=E_{min}$, then $\rho=\rho_{min}=\left(L^{2}/4c\right)^{1/6}=\mathrm{const}$, and
$\theta(\tau)=\frac{L}{\rho_{min}^{2}}\tau+C_{\theta}.$ (77)
@@
The general solution for the CCM with this potential can be found by Eqs. (52) and (58).
@@ VI CONCLUSIONS
In this paper, we have constructed new integrable chiral cosmological models with two-fields and potentials in terms of hyperbolic functions. These potentials are presented in Eqs. (60), (67), and (78). The general solutions have been found in terms of the Jacobi elliptic functions of the conformal time.
To get these solutions we have used the corresponding models in the Jordan frame, for which the Ricci scalar is an integral of motion. The integration of the initial system of equations reduces to the integration of system (37) of two second-order equations, that has one integral of motion for any potential $V$ given by Eq. (11). In this paper, we present examples of the potential $V$, for which system (37) has two independent integrals of motion and, hence, is integrable.
Two-field CCMs actively investigated in cosmology to describe the evolution of the observable Universe, including inflation, primordial black hole formation and dark energy [42, 43, 44, 45, 46, 47, 48, 49]. The constructed models can be considered as two-field generalizations of the single-field integrable model proposed in Ref. [5], which has been actively investigated as a possible description of the evolution of the early Universe. We hope that the proposed integrable models will be useful to describe the Universe evolution.

### v3

@@ V.1 The way to get the general cosmological solution of the CCM
In the FLRW metric (59), Eqs. (22)–(24) have the following form:
@@
Our goal is to get the general solutions of these equations for the potentials given by formula (21). If we know the general solutions for the corresponding model in the Jordan frame, or, in other words, if the functions $a(\tau)$, $\sigma^{1}(\tau)$, and $\sigma^{2}(\tau)$ are known, then the scale function $a_{\mathrm{E}}$ is given by Eq. (60) and the Hubble parameter
@@
Here we use the fact that conformal times for both frames coincide (up to an additive constant, which can be set to be zero without loss of generality), since $dt/a=dt_{\mathrm{E}}/a_{\mathrm{E}}$.
@@
V.2 Construction of integrable models
To get an integrable CCM we choose a particular form of the potential $\tilde{V}$ in the Jordan frame, for which system (37) is integrable. This system is a Hamiltonian one and has an integral of motion (39). To prove the integrability of a such Hamiltonian model one should get the second independent integral of motion or show that the model passes the Painleve test (the description of the Painleve test and the use of it in cosmology see, for example, Refs. [50, 51, 52, 53, 54, 55, 56]). Integrable models with polynomial potentials $\tilde{V}$ have been found in Refs. [51, 57]. Moreover, the full list of polynomial potentials, for which the second invariant is a polynomial in momenta of order four or less has been found in Ref. [57]. This list consists of two sets of potentials $\tilde{V}$. The first set includes potentials for which system (37) is integrable for $K=0$ only. These potentials are
$\tilde{V}_{a}=V_{0}\left[\frac{4}{3}\left(y^{1}\right)^{4}+\left(y^{1}\right)^{2}\left(y^{2}\right)^{2}+\frac{1}{6}\left(y^{2}\right)^{4}\right],$ (67)
$\tilde{V}_{b}=V_{0}\left[\frac{4}{3}\left(y^{1}\right)^{4}+\left(y^{1}\right)^{2}\left(y^{2}\right)^{2}+\frac{1}{12}\left(y^{2}\right)^{4}\right],$ (68)
and the potentials obtained by reflection: $y^{1}\rightarrow y^{2},\quad y^{2}\rightarrow y^{1}$. The corresponding second integrals are presented in Refs. [51, 57].
The second set corresponds to integrable systems with an arbitrary $K$. We consider these potentials in the next subsections in detail.
V.3 Potential $V=\frac{1}{2}{R_{0}U_{0}}$
@@
$V_{\mathrm{E}}=\frac{M^{4}_{\mathrm{Pl}}R_{0}}{8U_{0}}\cosh^{4}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\cosh^{4}\left(\frac{\chi^{2}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\,.$ (70)
@@
$y^{A}=\left\{\begin{aligned} B_{1}^{A}\sin(\sqrt{K}\tau)+B_{0}^{A}\cos(\sqrt{K}\tau),\quad\qquad&\quad K>0;\\ B_{1}^{A}\tau+B_{0}^{A},\qquad\qquad\qquad\quad\quad\quad\qquad\,\,\,&\quad K=0;\\ B_{1}^{A}\sinh(\sqrt{-K}\tau)+B^{A}_{0}\cosh(\sqrt{-K}\tau),\,\,&\quad K<0;\\ \end{aligned}\right.$ (71)
@@
V.4 Potential $V=\frac{1}{2}{R_{0}U_{0}}+c_{1}\left(\sigma^{1}\right)^{4}+c_{2}\left(\sigma^{2}\right)^{4}$
@@
System (37) with the potential (72), is completely separable, and we have two independent integrals of motion:
@@
System (73) includes two independent equations. Similar equations arise in the single-field model with hyperbolic potential, for which exact solutions with nonzero $K$ have been obtained in Ref. [9]. In the appendix of Ref. [9], the authors list the complete set of analytic solutions to these equations in the various regions of the parameter space.
In this paper, we write out only solutions of Eq. (73) with $K=0$; the form of these solutions depends on the signs of $c_{A}$ and $E_{A}$. Namely, we obtain:
@@
    $y^{A}(\tau)=\left(\frac{E_{A}}{c_{A}}\right)^{\frac{1}{4}}\mathrm{cn}\left.\left(\pm 2\left(E_{A}c_{A}\right)^{\frac{1}{4}}(\tau-\tau_{i})+\mathrm{cn}^{-1}\left(u_{A}\left|\frac{1}{2}\right.\right)\right|\frac{1}{2}\right);$ (74)
@@
Here constants $u_{A}=|c_{A}/E_{A}|^{1/4}y^{A}(\tau_{i})$ and the choice of sign in “$\pm$” are determined by initial conditions at $\tau=\tau_{i}$. Obviously, there is no solution when $E_{A}<0$ and $c_{A}>0$.
@@
If we put $\chi^{2}=0$, then $V_{\mathrm{E}}$ coincides with the potential from the single-field model proposed by Bars and Chen in Ref. [7]:
@@
So, the proposed integrable model with potential (77) is a two-field generalization of the known single-field integrable model [7, 9].
V.5 Generalization of the potential $V=\frac{1}{2}{R_{0}U_{0}}+c_{1}\left(\sigma^{1}\right)^{4}+c_{2}\left(\sigma^{2}\right)^{4}$
The potential $V_{1}$ can be generalized. Indeed, let us consider the following potential:
$\begin{split}\tilde{V}_{1m}&=c_{1}\left(\cos(\theta_{0})y^{1}+\sin(\theta_{0})y^{2}\right)^{4}+c_{2}\left(\cos(\theta_{0})y^{2}-\sin(\theta_{0})y^{1}\right)^{4}\\ &=\left(c_{1}\sin^{4}(\theta_{0})+c_{2}\cos^{4}(\theta_{0})\right)\left(y^{2}\right)^{4}+4\cos(\theta_{0})\sin(\theta_{0})\left(c_{1}\sin^{2}(\theta_{0})-c_{2}\cos^{2}(\theta_{0})\right){y^{1}}\left(y^{2}\right)^{3}\\ &{}+6\cos^{2}(\theta_{0})\sin^{2}(\theta_{0})\left(c_{1}+c_{2}\right)\left(y^{1}\right)^{2}\left(y^{2}\right)^{2}\\ &{}+4\cos(\theta_{0})\sin(\theta_{0})\left(c_{1}\cos^{2}(\theta_{0})-c_{2}\sin^{2}(\theta_{0})\right)\left(y^{1}\right)^{3}y^{2}+\left(c_{1}\cos^{4}(\theta_{0})+c_{2}\sin^{4}(\theta_{0})\right)\left(y^{1}\right)^{4},\end{split}$ (79)
where $\theta_{0}$ is a constant.
After the rotation
$\begin{split}y^{1}&=\cos(\theta_{0})z^{1}(\tau)-\sin(\theta_{0})z^{2}(\tau),\\ y^{2}&=\sin(\theta_{0})z^{1}(\tau)+\cos(\theta_{0})z^{2}(\tau),\end{split}$ (80)
we get
$\tilde{V}_{1}(z^{1},z^{2})=c_{1}\left(z^{1}\right)^{4}+c_{2}\left(z^{2}\right)^{4}.$ (81)
Note that the kinetic part of the Hamiltonian (38) and the quadratic part of the potential are invariant under transformation (80), and so we obtain that the model with the potential
$V_{1m}(\sigma^{1},\sigma^{2})=\tilde{V}_{1m}(\sigma^{1},\sigma^{2})+\frac{1}{2}R_{0}U_{0}$
is integrable.
Using Eq. (17), we get the potential in the Einstein frame:
$\begin{split}V_{E}&=36M^{4}_{\mathrm{Pl}}\cosh^{4}\left(\frac{\chi^{1}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\left[\frac{R_{0}}{288\,U_{0}}\cosh^{4}\left(\frac{\chi^{2}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right.\\ &\left.{}+\left[c_{1}\cos^{4}(\theta_{0})+c_{2}\sin^{4}(\theta_{0})\right]\sinh^{4}\left(\frac{\chi^{2}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right.\\ &\left.{}+4\cos(\theta_{0})\sin(\theta_{0})\left[c_{1}\cos^{2}(\theta_{0})-c_{2}\sin^{2}(\theta_{0})\right]\tanh\left(\frac{\chi^{1}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\sinh^{3}\left(\frac{\chi^{2}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right.\\ &\left.{}+6\left(c_{1}+c_{2}\right)\cos^{2}(\theta_{0})\sin^{2}(\theta_{0})\tanh^{2}\left(\frac{\chi^{1}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\sinh^{2}\left(\frac{\chi^{2}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right.\\ &\left.{}-4\cos(\theta_{0})\sin(\theta_{0})\left[c_{2}\cos^{2}(\theta_{0})-c_{1}\sin^{2}(\theta_{0})\right]\tanh^{3}\left(\frac{\chi^{1}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\sinh\left(\frac{\chi^{2}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right.\\ &\left.{}+\left[c_{1}\sin^{4}(\theta_{0})+c_{2}\cos^{4}(\theta_{0})\right]\tanh^{4}\left(\frac{\chi^{1}}{\sqrt{6}\,M_{\mathrm{Pl}}}\right)\right]\,.\end{split}$ (82)
The general solution for the CCM with this potential can be found by using Eqs. (60) and (66). Note that $\theta$, $c_{1}$, and $c_{2}$ are arbitrary parameters. So, we obtain new three-parametric sets of integrable models both in the Jordan and Einstein frames.
V.6 Potential $V=\frac{1}{2}{R_{0}U_{0}}+c\left(\left(\sigma^{1}\right)^{2}+\left(\sigma^{2}\right)^{2}\right)^{2}$
@@
$V_{2}=\frac{1}{2}{R_{0}U_{0}}+c\left(\left(\sigma^{1}\right)^{2}+\left(\sigma^{2}\right)^{2}\right)^{2},$ (83)
where $c$ is a nonzero constant, general solutions can also be explicitly written in terms of the Jacobi elliptic functions. In this case, Eq. (39) takes the form
$\frac{1}{2}\left(\frac{dy^{1}}{d\tau}\right)^{2}+\frac{1}{2}\left(\frac{dy^{2}}{d\tau}\right)^{2}+\frac{K}{2}\left(\left(y^{1}\right)^{2}+\left(y^{2}\right)^{2}\right)+c\left(\left(y^{1}\right)^{2}+\left(y^{2}\right)^{2}\right)^{2}=E.$ (84)
By introducing “polar coordinates,”
@@
$\frac{1}{2}\left(\frac{d\rho}{d\tau}\right)^{2}+\frac{1}{2}\rho^{2}\left(\frac{d\theta}{d\tau}\right)^{2}+\frac{K}{2}\rho^{2}+c\rho^{4}=E.$ (86)
In analogy with classical mechanics, one may observe that $\theta$ is a cyclic coordinate, which means that the second integral is
@@
$\frac{1}{2}\left(\frac{d\rho}{d\tau}\right)^{2}+\frac{L^{2}}{2\rho^{2}}+\frac{K}{2}\rho^{2}+c\rho^{4}=E.$ (88)
So, we obtain that the model with potential $V_{2}$ is integrable for an arbitrary value of $K$.
In what follows, we restrict ourselves to the case of $c>0$ only. In such a case the solution for $\rho$ reads
@@
$\rho^{2}_{0,1}=\alpha v_{0,1},\quad v_{k}=\cos\left(\frac{1}{3}\arccos\left(-\frac{\frac{K^{3}}{216}+\frac{1}{12}KEc+\frac{1}{4}L^{2}c^{2}}{\left(\frac{1}{3}Ec+\frac{1}{12}K^{2}\right)^{3/2}}\right)-\frac{2\pi k}{3}\right)-\frac{K}{2c\alpha},\quad k=0,1,2,$
$\alpha=2\sqrt{\frac{E}{3c}+\frac{K^{2}}{12c^{2}}},\quad\beta=\frac{1}{\sqrt{8c\alpha}}.$
The solution for $\theta$ is
$\theta(\tau)=\frac{2\beta L^{2}}{\alpha v_{0}\sqrt{v_{0}-v_{2}}}\Pi\left(\left.1-\frac{v_{1}}{v_{0}};\frac{\sqrt{v_{0}-v_{2}}}{2\beta}\tau+C_{\rho}\right|\frac{v_{0}-v_{1}}{v_{0}-v_{2}}\right)+C_{\theta}.$ (90)
@@
The general solution for the CCM with this potential can be found by using Eqs. (60) and (66).
@@ VI CONCLUSIONS
In this paper, we have constructed new integrable chiral cosmological models with two scalar fields and potentials in terms of hyperbolic functions. These potentials are presented in Eqs. (70), (77), (82), and (91). The general solutions have been found in terms of the Jacobi elliptic functions of the conformal time.
To get these solutions we have used the corresponding models in the Jordan frame, for which the Ricci scalar is an integral of motion. The integration of the initial system of equations reduces to the integration of system (37) of two second-order equations, that has one integral of motion for any potential $\tilde{V}$ given by Eq. (11). This system is a Hamiltonian one. The existence of nonzero spatial curvature adds quadratic terms proportional to $K$ to the Hamiltonian (38). In contrast to the single-field model, the existence of one integral of motion is not enough to obtain integrable models in the FLRW metric. In this paper, we construct new examples of the potential $\tilde{V}$, for which system (37) has two independent integrals of motion for an arbitrary $K$ and, hence, the corresponding two-field CCM is integrable not only in the spatially flat FLRW metric, but also in the open and closed FLRW metric. All integrable cases of the Hamiltonian (38) with fourth-order polynomial potentials, for which the second invariant is a polynomial in momenta of order four or less, have been found in Ref. [57]. To the best of our knowledge the existence of an integrable system (37) with a more complicated potential, for example, a nonpolynomial one, is an open question.
One of the important questions of the theoretical cosmology has to do with a possible evolution of the Universe before inflation. In particular, the question of the existence of a smooth connection between gravity and antigravity [47, 58, 59, 60] and the consideration of a cyclic universe [61, 62, 46, 63, 64] attract a lot of attention. Single-field integrable cosmological models are actively used in such investigations. Moreover, integrable models can be used as main parts for more realistic models of gravity that confirm with observational data. For example, the model proposed in Ref. [47] can describe the Universe before and at the beginning of inflation, and so exact solutions can be used to calculate inflationary parameters analytically, without resorting to the slow-roll approximation. At the same time, some new terms should be added to get the exit from inflation. A similar situation arises with the integrable model proposed in Ref. [13], with the goal to describe the possibility of producing a bouncing universe in the framework of scalar-tensor gravity. In this model, the Hubble parameter is a monotonically increasing function, and to get more realistic models with a nonmonotonic Hubble parameter one needs to slightly modify the specified integrable model [65, 66].
Two-field CCMs are actively used in cosmology to describe the evolution of the observable Universe, including inflation, primordial black hole formation and dark energy [5, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]. The constructed models can be considered as two-field generalizations of the single-field integrable model proposed in Ref. [7], which has been actively investigated as a possible description of the evolution of the early Universe. We hope that the proposed two-field integrable models will be useful for describing the Universe evolution.